OpenStudy (anonymous):

x2 - 12x + 27 please show me the steps please and the answer :) please and thank you very much

OpenStudy (amistre64):

the steps tooooo ....... what?

OpenStudy (anonymous):

factor by the way

OpenStudy (nenadmatematika):

=(x-9)(x-3)

OpenStudy (amistre64):

many ways to factor

OpenStudy (anonymous):

x^2 -12x+27

OpenStudy (amistre64):

+27 needs -- or ++ -12 says we need the -- part 9*3 = 27 9+3 = 12 (x-9)(x-3)

OpenStudy (nenadmatematika):

omg copy paste....:D